OpenStudy (anonymous):

1-cotx cscx-cot^2x How to simplified?

OpenStudy (anonymous):

\[1-\frac{ \cos x }{ \sin x }*\frac{ 1 }{ \sin x }-\frac{ \cos ^2x }{ \sin ^2x }=1-\frac{ \cos x-\cos ^2x }{ \sin ^2x }\]